OpenStudy (anonymous):

4x-2y=32 -3x-5y=-11. i need help fast.

OpenStudy (anonymous):

What method do you want to use?

OpenStudy (anonymous):

elimination using adition or subtraction

OpenStudy (anonymous):

Ok which one do you want eliminate first the x or the y?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

Ok so you multiply the first equation by -3 then the second one by 4

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

-12x-6y=32 -12x-20y=-11

OpenStudy (anonymous):

?

OpenStudy (anonymous):

it should be +6y on the top You also forgot to multiply the last number of each equation.

OpenStudy (anonymous):

=-96 =-44

OpenStudy (anonymous):

Yup. -12x+6y=-96 -12x-20y=-44 Now subtract the first line to the second one. What do you get?

OpenStudy (anonymous):

x+14y=150

OpenStudy (anonymous):

-12x+12x -20y -6x=-44+96 If you subtract a negative it turns into a postive so now simplify it. You were close.

OpenStudy (anonymous):

x-14y=-140/14=10

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

0r x-26y=-140

OpenStudy (anonymous):

hello???!!